5:50; 12:288; 9:162; 7:?

Grade:12th Pass

3 Answers

Latika Leekha
9 years ago
Hello student,
The given ratios are
5:50
12:288
9:162
7:y (let)
Now we try to analyze the given ratios.
5: 5*5*2 (50 = 5.5.2)
12: 12*12*2 (288 = 12.12.2)
9: 9*9*2 (162 = 9.9.2)
Hence, seeing this we can write y as y = 7*7*2 = 98.
Hence, 98 is the required answer.
vishal banshiwal
5 years ago
The given ratios are
5: 50 :: 12: 288 :: 9: ?
1.  5: 5*5*2 = 50 
2.  12: 12*12*2  = 288 
3.   9: 9*9*2 = 162 
 
 
 
 
 
Atharv lomte
5 years ago
QUE=5:50; 12:288; 9:162; 7:?
  1. 5:50 = 52*2=50
  2. 12:288=122*2=288
  3. 9:162=92*2=162
  4. similarly, 7:? = 72*2=98
  5. So our answer is 98.
​​Answered by Atharv Balajirao Lomte.
Student class8